OpenStudy (anonymous):

I must solve for y: x=1/3yz

OpenStudy (anonymous):

y = 1/3xz

OpenStudy (anonymous):

no... x=1/3yz

OpenStudy (anonymous):

I know. if the question is x=1/(3yz), then solving for y yields y=1/(3xz). If the question is x= (1/3)*yz, then the answer is y= 3(x/z)

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

which of the 2 question forms is correct. Then I will show you.